# LimiX: Unleashing Structured-Data Modeling Capability for Generalist Intelligence

### v1

@@ 3.2 Mask Pattern Design
We employ a heterogeneous schedule that interleaves cell-wise, column-wise, and block masks, enabling the model to practise recovering both isolated entries and coordinated subsets of variables. Cell-wise masks refine local conditional predictions; column-wise masks force the model to treat an entire feature as missing and infer it from the remaining attributes; block masks target higher-order dependencies across semantically related groups (e.g., demographics with outcomes, laboratory panels with diagnoses). Masking rates are stratified by variable type, prevalence, and dispersion to avoid overfitting to common categories and to limit the influence of high-variance continuous features, and we exclude degenerate patterns that remove nearly all informative context. This diversified schedule provides broad coverage of conditional relationships and prevents the model from specializing to a narrow reconstruction regime, yielding a more faithful approximation of the joint distribution. In practice, we sample the mask ratio in [0.1, 0.4] for LimiX.
@@ 4 Pretraining Data Generation
The performance of foundation models largely depends on the diversity and quality of pretraining data. To obtain a well-generalized foundation model for tabular data, we generate synthetic datasets using Directed Acyclic Graphs (DAGs), enabling the creation of datasets with diverse characteristics. The data generation process consists of three stages: DAG generation, data sampling, and task adaptation. First, a DAG is constructed to represent complex causal dependencies among variables. Then, a subset of these variables is sampled to define a specific problem, allowing LimiX to develop causal reasoning capabilities. Finally, the sampled data is processed to align with various downstream tasks. Compared with the data generation methods used in current foundation models  (Hollmann et al., 2025; Qu et al., 2025), we adopt a hierarchical generation paradigm that establishes causal dependencies in a more controllable and interpretable manner, while the sampling strategies further enhance the ability to generate datasets with diverse solvability and characteristics.
@@ 4.1 DAG Generation based on Hierarchical SCMs
  • Multilayer perceptrons (MLPs): To determine the architecture of these neural networks, we uniformly sample properties such as the number of linear layers and their associated activation functions. When multiple layers are included, the MLP introduces complex nonlinear dependencies along the edge, whereas with only a single layer, it degenerates into a simple transformation. For weight initialization, we randomly choose from Xavier initialization (Glorot & Bengio, 2010) and He initialization (He et al., 2015). The activation functions are uniformly sampled from identity mapping, hyperbolic tangent, sigmoid, logarithm, absolute value, sine, squaring, modulo operation, heaviside step function, and GELU.
@@ 5 Retrieval-based Ensemble
We adopt an inference-time and retrieval-based ensemble strategy that leverages LimiX's learned attention scores to upweight and select representative in-context samples and features without any additional training, so that we can further improve the performance of LimiX.
@@
In terms of retrieval, we perform two forward passes of LimiX for each pipeline. The first pass is performed based on all in-context samples and is employed for retrieval. The second pass is performed based on the customized in-context samples retrieved in the first pass.
@@
(a) Visualization of synthetic samples. Each sector represents a category of in-context samples. The yellow pentagram represents the test sample.
@@
Figure 5: Toy example of sample-level attention. In-context samples that share the same category as the query sample are assigned higher scores through the attention module of LimiX.
@@
Figure 6: Toy example of feature-level attention for the outcome variable. Direct causes of the outcome and the outcome itself are assigned almost all the attention weights in LimiX.
@@
  • Sample-level retrieval. We generate 2D synthetic data of 10 classes, where each class of data points occupies a sector of a circle in Figure 5(a). Since attention depicts sample similarity in the latent embedding space and it also takes the dependency between input features and the category label into account, it is capable of capturing more complex dependencies than naively using Euclidean distance in the original 2D space of features. From Figure 5(b), we can see that the attention module predominantly assigns large weights to in-context instances sharing the same category label as the query instance. This indicates that the model leverages class-consistent contextual information from in-context samples to assist its prediction.
@@ 6 Theoretical Analysis on Context-Conditional Masked Modeling
Let $\Omega$ denote the space for each feature in the table. We represent the $m$ in-context samples, each with $d$ features, by the random matrix $\mathbf{X}^{\text{ct}}\in\Omega^{m\times d}$. We use the random vector $\mathbf{X}^{\text{te}}=(X^{\text{te}}_{1},\dots,X^{\text{te}}_{d})\in\Omega^{d}$ to denote the test sample111 For simplicity, we assume (i) all features share the same space $\Omega$, though the results generalize to distinct spaces; (ii) we use $m$ instead of $m_{ct}$ in Section 2; and (iii) there is only one test sample, i.e., $m_{te}=1$ in Section 2.. Meanwhile, we do not explicitly introduce a target label $y$ in this section for simplicity; instead, we treat it as a particular dimension of both the in-context and test samples. Consequently, one dimension of $\mathbf{X}^{\text{te}}$ is unknown to the model during the test phase.
@@ 6.2 The Choice of Mask Number
In this subsection, we demonstrate that the choice $k$ in Equation 1 has great impact on models' performances and we choose $k>1$ due to both sample efficiency and generalization considerations. Our result is an extension from the analysis by Li et al. (2024c) on the properties of masked sequence prediction.
@@ 7.1 Classification
For the quantitative evaluation of classification performance, we utilize multiple benchmarks, including TALENT-CLS (Liu et al., 2024), OpenML-CC18 (Bischl et al., 2017), PFN-CLS (Hollmann et al., 2025), TabZilla (McElfresh et al., 2023), and TabArena (Erickson et al., 2025). Among these benchmarks, the datasets containing more than 50,000 training samples (The number of testing samples is not constrained), 10,000 features, or 10 target categories were excluded. This selection process resulted in a final collection of 179 datasets from TALENT-CLS, 62 from OpenML-CC18, 29 from PFN-CLS, 27 from TabZilla, and 33 from TabArena.
Furthermore, we introduce Balanced Comprehensive Challenging Omni-domain (BCCO) Benchmark, comprising BCCO-CLS and BCCO-REG, for the evaluation of LimiX and baseline models. The BCCO benchmark is constructed from extensive open-source structured-data corpora, meticulously deduplicated and cleaned. It presents a significant challenge due to several intrinsic characteristics: the distribution of dataset attributes (such as the ratio of categorical features), and the diversity of real-world prediction targets. Unlike previous benchmarks, nearly one-third of the datasets in our BCCO benchmark contain missing values.
The BCCO-CLS benchmark comprises 106 datasets. The collection spans diverse sources, domains, and scales, and is designed to cover a wide range of problem characteristics, including the number of samples, number of features, number of classes, categorical–to-numerical feature ratio, sample-to-feature ratio, and proportion of missing values. This diversity allows for uniform binning along these dimensions, enabling results to be reported within bins and macro-averaged across bins, thereby providing a nearly unbiased assessment of model performance across heterogeneous task regimes. Additionally, datasets containing more than 50,000 training samples (The number of testing samples is not constrained), 10,000 features, or 10 target categories were excluded.
@@
We compare LimiX with a range of state-of-the-art baseline models, categorized into tree-based models, neural networks (NN), and recent ICL-based approaches.
  • Tree-based approaches. We include XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), CatBoost (Dorogush et al., 2018), Random Forest (RF) (Breiman, 2001), and Extra Trees (ET) (Geurts et al., 2006). All models are optimized using the Optuna (Akiba et al., 2019) framework via 5-fold stratified cross-validation, with hyperparameters sampled from the ranges specified in Section A.2. Additionally, for AutoGluon-Tabular (Erickson et al., 2020), which automates workflows of model searching and ensemble, we use the default search space and set a default 600s time constraint for hyperparameter searching for each dataset.
  • NN-based approaches. We evaluate against SNN (Klambauer et al., 2017), AutoInt (Song et al., 2019), NODE (Popov et al., 2020), TabTransformer (Huang et al., 2020), GrowNet (Badirli et al., 2020), TabNet (Arik & Pfister, 2021), DCN-v2 (Wang et al., 2021), FT-Transformer (Gorishniy et al., 2021), MLP (Goodfellow et al., 2016; Gorishniy et al., 2021), ResNet (He et al., 2016; Gorishniy et al., 2021), SAINT (Somepalli et al., 2022), MLP-PLR (Gorishniy et al., 2022), DANets (Chen et al., 2022), TANGOS (Jeffares et al., 2023), T2G-Former (Yan et al., 2023), ExcelFormer (Chen et al., 2023b), Trompt (Chen et al., 2023c), TabR (Gorishniy et al., 2024), TabCaps (Chen et al., 2023a), RealMLP (Holzmüller et al., 2024), SwitchTab (Wu et al., 2024), and ModernNCA (Ye et al., 2025). These NN-based models are trained using the TALENT (Liu et al., 2024) Toolbox.
  • ICL-based models. Recent baselines includes TabPFN-v2 (Hollmann et al., 2025), TabICL (Qu et al., 2025), and Mitra (Zhang & Danielle, 2025).
Since TabDPT (Ma et al., 2024), another recent ICL-based model, is pretrained on real data that has a large overlap with datasets in the benchmarks, we do not adopt it as a baseline in the evaluation of standard classification and regression for a fair comparison. However, we include it in our experiments of fine-tuning in Section 7.6 to show the superiority of LimiX.
@@
The critical difference diagram is also employed to compare the performance differences between LimiX and the baseline models. We conducted a Friedman test followed by a post-hoc Wilcoxon-Holm test, using a significance level of 0.05. In the diagram, the horizontal line indicates the range of ranks among which differences are not statistically significant.
@@ Results.
Figure 7 shows that LimiX achieves the best performance among all methods on most datasets in BCCO-CLS. The most competitive baselines include other ICL-based models TabICL and TabPFN-v2, and the AutoML ensemble framework AutoGluon. In all critical difference diagrams, i.e. Figures 8, 9, 10, 11, 12 and 13, we can see that LimiX achieves the highest ranking in terms of all three evaluation metrics on BCCO-CLS, OpenML-CC18, and TALENT-CLS, where critical differences can be observed in most cases, indicating a significant margin.
For detailed quantitative results listed in Tables 1, 3, 5, 7, 9 and 11, LimiX outperforms all baselines on every benchmark in terms of both mean and rank of all three metrics. All these results clearly demonstrate that LimiX achieves state-of-the-art performance in terms of tabular classification tasks, surpassing not only traditional ensemble methods but also advanced ICL-based models.
@@
We use the sample subgroups when building BCCO-CLS to perform stratified analyses. Subgroups are defined based on the following criteria: the type of classification (binary or multi-class), the number of training samples, the ratio between the number of samples and features (length-to-width ratio), the proportion of categorical features, and the presence or absence of missing values. The number of training samples, categorical feature ratio, and length-to-width ratio are discretized into terciles (equal-frequency bins), ensuring that approximately the same number of datasets is allocated to each stratum and that the distribution of the dataset attribute which is not considered during analysis remains nearly uniform across strata.
From Figure 19, we observe that LimiX exhibits leading performance across all subgroups compared with other methods. Notably, after stratification, under some subgroups like the third subgroup in Figure 19(c) that indicates a larger training sample size, AutoGluon proves to be a strong competitor to ICL-based models. In these cases, LimiX is the only ICL-based model that outperforms AutoGluon. Notably, Figure 19(e) shows that, as the proportion of categorical features increases, performance for most baselines drops rapidly, reflecting the sparsity and high-cardinality challenges. In contrast, LimiX exhibits only modest degradation and remains comparatively stable across these regimes.
@@
BCCO-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.871±0.008 0.804±0.010 0.731±0.020 2.547 2.151 2.764 TabICL 0.847±0.008 0.768±0.010 0.672±0.016 6.642 8.142 10.132 AutoGluon 0.846±0.000 0.771±0.001 0.677±0.001 7.811 7.934 9.302 TabPFN-v2 0.843±0.005 0.772±0.007 0.679±0.009 8.462 9.094 10.698 Mitra 0.836±0.017 0.764±0.018 0.664±0.035 10.104 11.038 12.849 XGBoost 0.834±0.005 0.762±0.007 0.674±0.012 9.887 10.208 10.887 LightGBM 0.832±0.005 0.763±0.006 0.678±0.011 10.802 10.075 10.698 RF 0.829±0.007 0.756±0.008 0.652±0.014 11.349 11.717 13.047 CatBoost 0.829±0.008 0.757±0.010 0.664±0.014 11.943 11.000 11.792 ET 0.825±0.005 0.745±0.006 0.618±0.014 12.406 13.925 16.783 ModernNCA 0.815±0.017 0.752±0.016 0.658±0.027 15.377 15.689 15.321 FT-Transformer 0.813±0.014 0.744±0.015 0.642±0.030 13.509 13.953 14.377 ExcelFormer 0.810±0.018 0.742±0.015 0.655±0.023 15.783 15.217 13.840 TabR 0.809±0.019 0.750±0.019 0.657±0.035 16.443 14.415 13.443 T2G-Former 0.808±0.021 0.742±0.020 0.646±0.034 14.500 14.491 13.670 MLP-PLR 0.804±0.007 0.733±0.011 0.635±0.018 15.717 16.179 15.377 ResNet 0.800±0.017 0.728±0.016 0.641±0.022 16.943 17.160 15.896 TANGOS 0.799±0.015 0.731±0.016 0.641±0.022 17.179 16.943 15.406 DCN-v2 0.794±0.020 0.725±0.018 0.618±0.035 16.953 16.925 16.698 RealMLP 0.794±0.020 0.738±0.014 0.644±0.021 18.255 16.462 15.208 SAINT 0.791±0.017 0.726±0.013 0.623±0.021 18.736 17.642 16.311 MLP 0.787±0.011 0.720±0.013 0.614±0.018 18.509 17.925 18.783 AutoInt 0.779±0.028 0.718±0.017 0.601±0.037 20.208 19.453 19.849 SNN 0.773±0.024 0.708±0.019 0.584±0.049 20.698 20.321 20.594 DANets 0.771±0.029 0.705±0.027 0.601±0.043 20.377 19.575 19.868 SwitchTab 0.766±0.021 0.700±0.018 0.590±0.036 21.396 21.991 20.868 TabTransformer 0.762±0.018 0.699±0.023 0.566±0.047 20.670 19.849 20.321 NODE 0.754±0.046 0.695±0.031 0.531±0.078 23.925 22.594 25.255 TabCaps 0.744±0.027 0.701±0.019 0.580±0.026 23.708 21.670 22.113 TabNet 0.712±0.047 0.685±0.028 0.561±0.041 27.151 24.660 24.774 GrowNet 0.682±0.033 0.641±0.028 0.522±0.040 27.726 27.368 26.387
@@
TALENT-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.903±0.004 0.861±0.005 0.752±0.016 3.894 3.000 4.028 TabPFN-v2 0.895±0.002 0.850±0.003 0.727±0.008 6.056 5.860 7.670 TabICL 0.894±0.003 0.845±0.005 0.715±0.008 5.536 6.480 7.955 AutoGluon 0.891±0.000 0.845±0.000 0.719±0.001 6.229 6.642 7.425 Mitra 0.882±0.000 0.834±0.001 0.689±0.001 10.391 11.095 13.168 XGBoost 0.881±0.003 0.837±0.004 0.713±0.009 9.235 9.380 9.715 LightGBM 0.880±0.003 0.836±0.004 0.713±0.009 9.486 9.721 10.067 RF 0.877±0.003 0.828±0.004 0.691±0.010 12.307 12.793 13.927 CatBoost 0.876±0.005 0.828±0.006 0.704±0.014 11.453 11.497 11.743 ET 0.875±0.003 0.821±0.006 0.662±0.017 12.994 15.380 16.933 ExcelFormer 0.870±0.012 0.826±0.009 0.699±0.019 13.358 13.989 12.961 ResNet 0.866±0.009 0.825±0.006 0.695±0.014 14.933 14.721 13.207 MLP 0.862±0.008 0.817±0.008 0.675±0.017 16.419 16.140 16.346 TANGOS 0.861±0.009 0.818±0.007 0.684±0.020 16.011 15.978 14.754 ModernNCA 0.861±0.013 0.825±0.010 0.683±0.022 17.101 16.877 16.151 FT-Transformer 0.859±0.010 0.822±0.007 0.678±0.017 15.693 15.905 15.609 TabR 0.858±0.015 0.824±0.011 0.680±0.026 17.413 16.034 15.179 T2G-Former 0.858±0.016 0.823±0.010 0.683±0.019 15.777 15.536 15.134 DCN-v2 0.854±0.016 0.815±0.013 0.662±0.027 17.849 18.039 17.899 MLP-PLR 0.849±0.007 0.816±0.009 0.663±0.019 18.492 17.346 17.391 DANets 0.848±0.018 0.805±0.018 0.654±0.038 19.039 18.838 18.497 AutoInt 0.842±0.011 0.803±0.008 0.646±0.022 20.587 21.246 20.721 SwitchTab 0.842±0.013 0.795±0.011 0.637±0.029 21.346 21.894 21.089 RealMLP 0.839±0.017 0.820±0.007 0.678±0.017 20.212 17.687 15.994 SNN 0.836±0.020 0.796±0.016 0.625±0.041 22.011 21.955 22.089 TabCaps 0.834±0.014 0.813±0.008 0.654±0.017 21.324 18.223 18.687 TabTransformer 0.832±0.014 0.790±0.016 0.627±0.037 21.743 21.816 20.933 NODE 0.830±0.043 0.779±0.040 0.570±0.104 22.883 23.240 24.816 TabNet 0.818±0.035 0.794±0.021 0.630±0.041 25.006 22.603 23.039 SAINT 0.813±0.012 0.781±0.007 0.630±0.020 20.279 19.402 18.888 GrowNet 0.743±0.019 0.704±0.011 0.542±0.024 27.453 27.000 25.028

### v2

@@ 3.2 Mask Pattern Design
We employ a heterogeneous schedule that interleaves cell-wise, column-wise, and block masks, enabling the model to practise recovering both isolated entries and coordinated subsets of variables. Cell-wise masks refine local conditional predictions; column-wise masks force the model to treat an entire feature as missing and infer it from the remaining attributes; block masks target higher-order dependencies across semantically related groups (e.g., demographics with outcomes, laboratory panels with diagnoses). Masking rates are stratified by variable type, prevalence, and dispersion to avoid overfitting to common categories and to limit the influence of high-variance continuous features, and we exclude degenerate patterns that remove nearly all informative context. This diversified schedule provides broad coverage of conditional relationships and prevents the model from specializing to a narrow reconstruction regime, yielding a more faithful approximation of the joint distribution. In practice, we sample the mask ratio in [0.1, 0.4] for LimiX models.
@@ 4 Pretraining Data Generation
Performance of foundation models largely depends on the diversity and quality of pretraining data. To obtain a well-generalized foundation model for tabular data, we generate synthetic datasets using Directed Acyclic Graphs (DAGs), enabling the creation of datasets with diverse characteristics. The data generation process consists of three stages: DAG generation, data sampling, and task adaptation. First, a DAG is constructed to represent complex causal dependencies among variables. Then, a subset of these variables is sampled to define a specific problem, allowing LimiX models to develop causal reasoning capabilities. Finally, the sampled data is processed to align with various downstream tasks. Compared with the data generation methods used in current foundation models (hollmann2025tabpfn_nature; qu2025tabicl), we adopt a hierarchical generation paradigm that establishes causal dependencies in a more controllable and interpretable manner, while the sampling strategies further enhance the ability to generate datasets with diverse solvability and characteristics.
@@ 4.1 DAG Generation based on Hierarchical SCMs
  • Multilayer perceptrons (MLPs): To determine the architecture of these neural networks, we uniformly sample properties such as the number of linear layers and their associated activation functions. When multiple layers are included, the MLP introduces complex nonlinear dependencies along the edge, whereas with only a single layer, it degenerates into a simple transformation. For weight initialization, we randomly choose from Xavier initialization (glorot2010understanding) and He initialization (he2015delving). The activation functions are uniformly sampled from identity mapping, hyperbolic tangent, sigmoid, logarithm, absolute value, sine, squaring, modulo operation, heaviside step function, and GELU.
@@ 5 Retrieval-based Ensemble
We adopt an inference-time and retrieval-based ensemble strategy that leverages LimiX's learned attention scores to upweight and select representative in-context samples and features without any additional training, so that we can further improve the performance of LimiX models.
@@
In terms of retrieval, we perform two forward passes of LimiX models for each pipeline. The first pass is performed based on all in-context samples and is employed for retrieval. The second pass is performed based on the customized in-context samples retrieved in the first pass.
@@
(a) Visualization of synthetic samples. Each sector represents a category of in-context samples. The yellow pentagram represent the test sample.
@@
Figure 5: Toy example of sample-level attention. In-context samples that share the same category as the query sample are assigned higher scores through the attention module of LimiX-16M.
@@
Figure 6: Toy example of feature-level attention for the outcome variable. Direct causes of the outcome and the outcome itself are assigned almost all the attention weights in LimiX-16M.
@@
  • Sample-level retrieval. We generate 2D synthetic data of 10 classes, where each class of data points occupy a sector of a circle in Figure 5(a).
    Since attention depicts sample similarity in the latent embedding space and it also takes the dependency between input features and the category label into account, it is capable of capturing more complex dependencies than naively using Euclidean distance in the original 2D space of features. From Figure 5(b), we can see that the attention module predominantly assigns large weights to in-context instances sharing the same category label as the query instance. This indicates that the model leverages class-consistent contextual information from in-context samples to assist its prediction.
@@ 6 Theoretical Analysis on Context-Conditional Masked Modeling
Let $\Omega$ denote the space for each feature in the table. We represent the $m$ in-context samples, each with $d$ features, by the random matrix $\mathbf{X}^{\text{ct}}\in\Omega^{m\times d}$. We use the random vector $\mathbf{X}^{\text{te}}=(X^{\text{te}}_{1},\dots,X^{\text{te}}_{d})\in\Omega^{d}$ to denote the test sample [^1] For simplicity, we assume (i) all features share the same space $\Omega$, though the results generalize to distinct spaces; (ii) we use $m$ instead of $m_{ct}$ in Section 2; and (iii) there is only one test sample, i.e., $m_{te}=1$ in Section 2.. Meanwhile, we do not explicitly introduce a target label $y$ in this section for simplicity; instead, we treat it as a particular dimension of both the in-context and test samples. Consequently, one dimension of $\mathbf{X}^{\text{te}}$ is unknown to the model during the test phase.
@@ 6.2 The Choice of Mask Number
In this subsection, we demonstrate that the choice $k$ in Equation 1 has great impact on models' performances and we choose $k>1$ due to both sample efficiency and generalization considerations. Our result is an extension from the analysis by li2024promises on the properties of masked sequence prediction.
@@ 7.1 Classification
For the quantitative evaluation of classification performance, we utilize multiple benchmarks, including TALENT-CLS (liu2024talenttabularanalyticslearning), OpenML-CC18 (bischl2017openml), PFN-CLS (hollmann2025tabpfn_nature), TabZilla (mcelfresh2023neuralzilla), and TabArena-CLS (erickson2025tabarena). Among these benchmarks, the datasets containing more than 50,000 training samples (The number of testing samples is not constrained), 10,000 features, or 10 target categories were excluded. This selection process resulted in a final collection of 179 datasets from TALENT-CLS, 62 from OpenML-CC18, 29 from PFN-CLS, 27 from TabZilla, and 33 from TabArena-CLS.
Furthermore, we introduce Balanced Comprehensive Challenging Omni-domain (BCCO) Benchmark, comprising BCCO-CLS and BCCO-REG, for the evaluation of LimiX models and baseline models. The BCCO benchmark is constructed from extensive open-source structured-data corpora, meticulously deduplicated and cleaned. It presents a significant challenge due to several intrinsic characteristics: the distribution of dataset attributes (such as the ratio of categorical features), and the diversity of real-world prediction targets. Unlike previous benchmarks, nearly one-third of the datasets in our BCCO benchmark contain missing values. The BCCO benchmarks are publicly available at https://huggingface.co/datasets/stableai-org/bcco_cls and https://huggingface.co/datasets/stableai-org/bcco_reg.
The BCCO-CLS benchmark comprising 106 datasets. The collection spans diverse sources, domains, and scales, and is designed to cover a wide range of problem characteristics, including the number of samples, number of features, number of classes, categorical–to-numerical feature ratio, sample-to-feature ratio, and proportion of missing values. This diversity allows for uniform binning along these dimensions, enabling results to be reported within bins and macro-averaged across bins, thereby providing a nearly unbiased assessment of model performance across heterogeneous task regimes. Additionally, datasets containing more than 50,000 training samples (The number of testing samples is not constrained), 10,000 features, or 10 target categories were excluded.
@@
We compare LimiX models with a range of state-of-the-art baseline models, categorized into tree-based models, neural networks (NN), and recent ICL-based approaches.
  • Tree-based approaches. We include XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), CatBoost (dorogush2018catboost), Random Forest (RF) (breiman2001random), and Extra Trees (ET) (geurts2006extremely). All models are optimized using the Optuna (akiba2019optuna) framework via 5-fold stratified cross-validation, with hyperparameters sampled from the ranges specified in Section A.2. Additionally, for AutoGluon-Tabular (erickson2020autogluon), which automates workflows of model searching and ensemble, we use the default search space and set a default 600s time constraint for hyperparameter searching for each dataset.
  • NN-based approaches. We evaluate against SNN (klambauer2017self), AutoInt (song2019autoint), NODE (popov2019neuralnode), TabTransformer (huang2020tabtransformer), GrowNet (badirli2020gradient), TabNet (arik2021tabnet), DCN-v2 (wang2021dcn), FT-Transformer (gorishniy2021revisiting), MLP (goodfellow2016deep; gorishniy2021revisiting), ResNet (he2016deepresnet; gorishniy2021revisiting), SAINT (somepalli2021saint), MLP-PLR (gorishniy2022embeddingsmlpplr), DANets (chen2022danets), TANGOS (jeffares2023tangos), T2G-Former (yan2023t2g), ExcelFormer (chen2023excelformer), Trompt (chen2023trompt), TabR (gorishniy2023tabr), TabCaps (chen2023tabcaps), RealMLP (holzmuller2024betterrealmlp), SwitchTab (wu2024switchtab), and ModernNCA (ye2024revisitingmodernnca). These NN-based models are trained using the TALENT (liu2024talenttabularanalyticslearning) Toolbox.
  • ICL-based models. Recent baselines includes TabPFN-v2 (hollmann2025tabpfn_nature), TabICL (qu2025tabicl), and Mitra (exampleWebsite).
Since TabDPT (ma2024tabdpt), another recent ICL-based model, is pretrained on real data that has a large overlap with datasets in the benchmarks, we do not adopt it as a baseline in the evaluation of standard classification and regression for a fair comparison. However, we include it in our experiments of fine-tuning in Section 7.6 to show the superiority of LimiX models.
@@
The critical difference diagram is also employed to compare the performance differences between LimiX models and the baseline models. We conducted a Friedman test followed by a post-hoc Wilcoxon-Holm test, using a significance level of 0.05. In the diagram, the horizontal line indicates the range of ranks among which differences are not statistically significant.
@@ Results.
Figure 7 shows that LimiX-16M achieves the best performance among all methods on most datasets in BCCO-CLS, followed closely by LimiX-2M. The most competitive baselines include other ICL-based models TabICL and TabPFN-v2, and the AutoML ensemble framework AutoGluon. In all critical difference diagrams, i.e. Figures 8, 9, 10, 11, 12 and 13, we can see that LimiX-16M achieves the highest ranking in terms of all three evaluation metrics on BCCO-CLS, OpenML-CC18, and TALENT-CLS, where critical differences can be observed in most cases, indicating a significant margin. Despite operating under limited computational resources, LimiX-2M still delivers highly competitive results.
For detailed quantitative results listed in Tables 1, 3, 5, 7, 9 and 11, LimiX-16M outperforms all baselines on every benchmark in terms of both mean and rank of all three metrics. All these results clearly demonstrate that LimiX-16M achieves state-of-the-art performance in terms of tabular classification tasks, surpassing not only traditional ensemble methods but also advanced ICL-based models. Across all classification benchmarks, LimiX-2M also outperforms other strong baselines, despite operating under highly limited computational resources.
@@
We use the sample subgroups when building BCCO-CLS to perform stratified analyses. Subgroups are defined based on the following criteria: the type of classification (binary or multi-class), the number of training samples, the ratio between the number of samples and features (length-to-width ratio), the proportion of categorical features, and the presence or absence of missing values. The number of training samples, categorical feature ratio, and length-to-width ratio ared discretized into terciles (equal-frequency bins), ensuring that approximately the same number of datasets is allocated to each stratum and that the distribution of the dataset attribute which is not considered during analysis remains nearly uniform across strata.
From Figure 20, we observe that LimiX-16M exhibits leading performance across all subgroups compared with other methods. Notably, after stratification, under some subgroups like the third subgroup in Figure 20(c) that indicates a larger training sample size, AutoGluon proves to be a strong competitor to ICL-based models. In these cases, LimiX-16M is the only ICL-based model that outperforms AutoGluon. Notably, Figure 20(e) shows that, as the proportion of categorical features increases, performance for most baselines drops rapidly, reflecting the sparsity and high-cardinality challenges. In contrast, LimiX-16M exhibits only modest degradation and remains comparatively stable across these regimes.
@@
BCCO-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX-16M 0.871 0.804 0.731 2.642 2.245 2.887 LimiX-2M 0.858 0.787 0.701 6.170 6.481 8.538 TabICL 0.847 0.768 0.672 7.255 8.726 10.783 AutoGluon 0.846 0.771 0.677 8.500 8.500 9.934 TabPFN-v2 0.843 0.772 0.679 9.208 9.821 11.396 XGBoost 0.834 0.762 0.674 10.660 10.849 11.547 Mitra 0.836 0.764 0.664 10.953 11.698 13.585 LightGBM 0.832 0.763 0.678 11.594 10.811 11.406 RF 0.829 0.756 0.652 12.170 12.406 13.726 CatBoost 0.829 0.757 0.664 12.792 11.755 12.557 ET 0.825 0.745 0.618 13.292 14.745 17.585 FT-Transformer 0.813 0.744 0.642 14.340 14.708 15.104 T2G-Former 0.808 0.742 0.646 15.321 15.170 14.377 ModernNCA 0.815 0.752 0.658 16.217 16.434 16.047 MLP-PLR 0.804 0.733 0.635 16.566 16.915 16.113 ExcelFormer 0.810 0.742 0.655 16.660 16.009 14.566 TabR 0.809 0.750 0.657 17.274 15.057 14.104 DCN-v2 0.794 0.725 0.618 17.783 17.736 17.453 ResNet 0.800 0.728 0.641 17.783 17.943 16.670 TANGOS 0.799 0.731 0.641 18.057 17.717 16.151 RealMLP 0.794 0.738 0.644 19.132 17.236 15.934 MLP 0.787 0.720 0.614 19.387 18.764 19.594 SAINT 0.791 0.726 0.623 19.604 18.415 17.066 AutoInt 0.779 0.718 0.601 21.113 20.292 20.660 DANets 0.771 0.705 0.601 21.302 20.377 20.708 TabTransformer 0.762 0.699 0.566 21.557 20.698 21.170 SNN 0.773 0.708 0.584 21.585 21.160 21.415 SwitchTab 0.766 0.700 0.590 22.340 22.849 21.764 TabCaps 0.744 0.701 0.580 24.670 22.632 23.019 NODE 0.754 0.695 0.531 24.877 23.443 26.198 TabNet 0.712 0.685 0.561 28.123 25.547 25.679 GrowNet 0.682 0.641 0.522 28.679 28.330 27.321
@@
TALENT-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX-16M 0.903 0.861 0.752 4.073 3.190 4.246 LimiX-2M 0.897 0.853 0.734 5.804 5.642 6.994 TabICL 0.894 0.845 0.715 6.073 7.028 8.514 TabPFN-v2 0.895 0.850 0.727 6.670 6.458 8.296 AutoGluon 0.891 0.845 0.719 6.899 7.291 8.067 XGBoost 0.881 0.837 0.713 10.034 10.123 10.425 LightGBM 0.880 0.836 0.713 10.307 10.475 10.777 Mitra 0.882 0.834 0.689 11.162 11.849 13.916 CatBoost 0.876 0.828 0.704 12.302 12.302 12.508 RF 0.877 0.828 0.691 13.179 13.637 14.754 ET 0.875 0.821 0.662 13.877 16.229 17.788 ExcelFormer 0.870 0.826 0.699 14.212 14.855 13.749 ResNet 0.866 0.825 0.695 15.793 15.514 13.955 FT-Transformer 0.859 0.822 0.678 16.536 16.715 16.413 T2G-Former 0.858 0.823 0.683 16.615 16.374 15.933 TANGOS 0.861 0.818 0.684 16.905 16.810 15.542 MLP 0.862 0.817 0.675 17.302 16.989 17.162 ModernNCA 0.861 0.825 0.683 17.955 17.693 16.966 TabR 0.858 0.824 0.680 18.263 16.804 15.922 DCN-v2 0.854 0.815 0.662 18.704 18.849 18.687 MLP-PLR 0.849 0.816 0.663 19.385 18.168 18.201 DANets 0.848 0.805 0.654 19.961 19.721 19.374 RealMLP 0.839 0.820 0.678 21.106 18.525 16.793 SAINT 0.813 0.781 0.630 21.162 20.274 19.715 AutoInt 0.842 0.803 0.646 21.492 22.156 21.587 TabCaps 0.834 0.813 0.654 22.257 19.106 19.570 SwitchTab 0.842 0.795 0.637 22.274 22.777 21.989 TabTransformer 0.832 0.790 0.627 22.637 22.598 21.760 SNN 0.836 0.796 0.625 22.933 22.855 22.972 NODE 0.830 0.779 0.570 23.810 24.145 25.726 TabNet 0.818 0.794 0.630 25.983 23.553 23.983 GrowNet 0.743 0.704 0.542 28.430 27.933 25.933